# Small Ramsey numbers for books, wheels, and generalizations

### v1

@@ Abstract
This brief note gives several new upper and lower bounds on Ramsey numbers for books and wheels, including tight lower bounds establishing the Ramsey numbers $R(B_{3},B_{6})=19$, $R(B_{4},B_{5})=19$, $R(B_{5},B_{6})=23$, and $R(B_{6},B_{7})=27$, and tight upper and lower bounds establishing $R(W_{5},W_{7})=15$. The new upper bound was obtained using the flag algebra method, and the lower bounds come from constructions discovered using tabu search.
@@ 1 Introduction
There are a number of results on Ramsey numbers for books and wheels, summarized in [rad_survey]. In particular, the first and third authors proved upper bounds for a variety of small Ramsey numbers, including for books and wheels, using the flag algebra method [LidP]; we establish the upper bounds here using the same method. A very recent paper [RL_bounds] used reinforcement learning to find Ramsey lower bound constructions for graphs including books and wheels. The idea of finding graph constructions with reinforcement learning was pioneered in [wagner], and it is an exciting and promising technique. But it currently seems to struggle to scale to larger problems in comparison with traditional local search methods [deepmind, RL2, RL1]. Indeed, here we improve each of the lower bounds for books and wheels from [RL_bounds] using tabu search.
@@ 2 Results
$R(W_{5},W_{7})$ 15* 15* 14 [RL_bounds] 16 [LidP]
$R(W_{5},W_{9})$ 18 17 [rad_survey]
$R(B_{2},B_{8})$ 21 19 [RS78] 22 [b2_uppers]
@@
$R(B_{2},B_{10})$ 24 23 [RS78] 26 [b2_uppers]
$R(B_{3},B_{6})$ 19* 17 [RL_bounds] 19 [LidP]
$R(B_{3},B_{7})$ 20
$R(B_{4},B_{5})$ 19* 18 [RL_bounds] 19 [RS78]
$R(B_{4},B_{7})$ 22
@@ 3 Method
Each of the lower bounds is established by a construction provided in Appendix A. The constructions were found using tabu search and verified independently in SageMath [sagemath]; the code for books and wheels is available at https://github.com/gwen-mckinley/ramsey-books-wheels. The upper bounds were found by semidefinite programming using the flag algebra method, as described in [LidP]. The certificates for the upper bounds are available at http://lidicky.name/pub/gr. Additionally, for $GR(3,K_{4},2)$ and $GR(4,K_{4},2)$, we were able to use brute force to enumerate all possible constructions for the lower bound. There are 263 such constructions for $GR(3,K_{4},2)$ and only one for $GR(4,K_{4},2)$.
3.1 Tabu search details
@@ 3.2 Scoring functions
3.3 Flag Algebras
The upper bounds were obtained using the flag algebra method. This method was developed by Razborov [MR2371204], and the main idea is to find a certificate for the bound as a sum of squares, which can be done efficiently using semidefinite programming. We used CSDP [csdp] and MOSEK [mosek] solvers in this work. While the solvers produce numerical results, we created tools to change the solution to an exact arithmetic one in SageMath [sagemath]. The flag algebra method is quite general, and it is naturally applicable to problems where one is interested in asymptotic results. For example, the Turán density of small hypergraphs and its variants [l2norm] is a group of problems where flag algebras led to many new improvements.
@@ 4 Further directions
4.1 Lower bounds
To extend the methods in this work for finding lower bounds, there are a number of natural directions to explore. Perhaps most obvious would be to implement everything in a lower-level language – all the code for wheels and books is currently written in Python. In particular, the score-change functions dominate the total cost, and speeding them up would almost certainly improve performance; relevant here is an extensive literature on subgraph counting (see [subgraph_survey]), and it may be possible to simply incorporate existing open-source solutions.
One could also experiment with other local search methods. Previous work on Ramsey lower bounds has used many small variations on tabu search and simulated annealing; some of these methods also achieve considerable speedup by incorporating assumptions about the structure of the final construction (e.g., symmetries, cyclic structure, or subgraph counts), for example, [whatsR55, exoo4_6, exoo13]. More broadly, there is a wide literature on local search algorithms [SLS_book], any of which could potentially be applied here. We experimented briefly with a number of possibilities, but did not find any that offered a clear improvement; nevertheless, it is very likely that an appropriate choice could significantly improve performance.
Last but not least, one could try to be be more strategic in choosing where to start a local search. One common approach is to build up from smaller graphs: we first find a construction on a smaller number of vertices avoiding “bad” subgraphs, then add a vertex, and run our local search starting from this new graph (as opposed to starting from a random graph) [whatsR55, exoo4_6]. The hope is that by doing this, we direct the search to a promising area of the (very large) search space. We also tried this method, with limited success, but it is likely worth further exploration.
@@ Acknowledgments
The second author would like to thank David Hemminger and Larry Wilson for many helpful and stimulating discussions, and for their consistent encouragement. This work used the computing resources at the Center for Computational Mathematics, University of Colorado Denver, including the Alderaan cluster, supported by the National Science Foundation award OAC-2019089.
@@
Appendix A Constructions
@@
    Adjacency matrix of a 14-vertex graph with no $W_{5}$ in color 0 and no $W_{7}$ in color 1:
    [[0,0,1,1,0,0,1,1,1,1,1,0,0,0],
     [0,0,1,0,0,0,1,1,1,0,1,1,0,0],
     [1,1,0,1,1,1,1,0,0,0,0,0,1,0],
     [1,0,1,0,0,1,1,0,0,0,0,1,1,1],
     [0,0,1,0,0,1,1,0,1,0,0,1,1,0],
     [0,0,1,1,1,0,0,0,0,1,0,1,0,1],
     [1,1,1,1,1,0,0,1,0,0,1,0,0,0],
     [1,1,0,0,0,0,1,0,1,1,0,0,0,1],
     [1,1,0,0,1,0,0,1,0,1,1,1,0,0],
     [1,0,0,0,0,1,0,1,1,0,1,0,1,1],
     [1,1,0,0,0,0,1,0,1,1,0,0,0,1],
     [0,1,0,1,1,1,0,0,1,0,0,0,1,1],
     [0,0,1,1,1,0,0,0,0,1,0,1,0,1],
     [0,0,0,1,0,1,0,1,0,1,1,1,1,0]]
@@
    Adjacency matrix of a 17-vertex graph with no $W_{5}$ in color 0 and no $W_{9}$ in color 1:
    [[0,0,0,0,0,0,1,1,1,0,0,1,1,1,0,0,1],
     [0,0,1,1,0,1,0,1,0,1,1,0,0,0,0,1,0],
     [0,1,0,0,1,1,0,0,0,1,1,0,1,1,0,1,0],
     [0,1,0,0,1,1,1,0,0,1,1,0,0,1,0,1,0],
     [0,0,1,1,0,1,0,0,1,1,1,1,0,0,0,1,1],
     [0,1,1,1,1,0,0,1,0,1,0,0,0,0,1,1,0],
     [1,0,0,1,0,0,0,1,1,0,0,1,0,1,1,0,1],
     [1,1,0,0,0,1,1,0,0,0,1,1,1,1,1,0,1],
     [1,0,0,0,1,0,1,0,0,1,0,1,1,0,1,1,1],
     [0,1,1,1,1,1,0,0,1,0,1,0,0,0,0,0,0],
     [0,1,1,1,1,0,0,1,0,1,0,0,0,0,1,1,0],
     [1,0,0,0,1,0,1,1,1,0,0,0,1,1,1,0,0],
     [1,0,1,0,0,0,0,1,1,0,0,1,0,1,1,0,1],
     [1,0,1,1,0,0,1,1,0,0,0,1,1,0,1,0,1],
     [0,0,0,0,0,1,1,1,1,0,1,1,1,1,0,0,1],
     [0,1,1,1,1,1,0,0,1,0,1,0,0,0,0,0,0],
     [1,0,0,0,1,0,1,1,1,0,0,0,1,1,1,0,0]]
@@
    Adjacency matrix of a 20-vertex graph with no $B_{2}$ in color 0 and no $B_{8}$ in color 1:
    [[0,1,0,1,0,1,0,1,1,0,0,1,1,0,1,1,1,0,0,1],
     [1,0,0,1,1,1,1,0,1,1,1,0,1,0,0,1,1,0,0,1],
     [0,0,0,0,0,0,1,1,0,1,1,1,1,1,1,0,1,1,1,1],
     [1,1,0,0,1,1,1,1,0,0,0,0,1,0,0,1,0,1,1,1],
     [0,1,0,1,0,1,1,0,1,1,1,0,0,1,0,1,0,1,1,0],
     [1,1,0,1,1,0,0,1,1,0,1,1,0,1,0,1,0,1,0,1],
     [0,1,1,1,1,0,0,0,0,1,1,0,1,0,0,0,1,1,1,1],
     [1,0,1,1,0,1,0,0,0,0,0,1,1,1,1,1,0,1,1,1],
     [1,1,0,0,1,1,0,0,0,1,1,1,0,1,1,1,1,0,0,0],
     [0,1,1,0,1,0,1,0,1,0,1,0,1,1,1,1,1,0,1,0],
     [0,1,1,0,1,1,1,0,1,1,0,1,0,1,0,0,1,1,0,1],
     [1,0,1,0,0,1,0,1,1,0,1,0,0,1,1,0,1,1,0,1],
     [1,1,1,1,0,0,1,1,0,1,0,0,0,0,1,1,1,0,1,1],
     [0,0,1,0,1,1,0,1,1,1,1,1,0,0,1,1,0,1,1,0],
     [1,0,1,0,0,0,0,1,1,1,0,1,1,1,0,1,1,0,1,0],
     [1,1,0,1,1,1,0,1,1,1,0,0,1,1,1,0,0,0,1,0],
     [1,1,1,0,0,0,1,0,1,1,1,1,1,0,1,0,0,0,0,1],
     [0,0,1,1,1,1,1,1,0,0,1,1,0,1,0,0,0,0,1,1],
     [0,0,1,1,1,0,1,1,0,1,0,0,1,1,1,1,0,1,0,0],
     [1,1,1,1,0,1,1,1,0,0,1,1,1,0,0,0,1,1,0,0]]
@@
    Adjacency matrix of a 21-vertex graph with no $B_{2}$ in color 0 and no $B_{9}$ in color 1:
    [[0,1,0,1,0,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0,1],
     [1,0,0,1,1,0,0,1,1,1,0,0,0,1,1,0,1,1,0,1,1],
     [0,0,0,0,0,1,0,1,0,0,1,1,1,1,0,1,1,1,1,1,1],
     [1,1,0,0,1,1,1,1,1,0,0,1,1,0,1,0,1,1,0,0,0],
     [0,1,0,1,0,0,1,0,1,1,0,0,1,0,1,1,1,1,1,1,0],
     [1,0,1,1,0,0,1,1,0,0,1,1,1,1,1,1,1,0,0,0,0],
     [1,0,0,1,1,1,0,0,1,1,1,1,1,0,1,1,0,0,1,0,0],
     [1,1,1,1,0,1,0,0,1,0,1,1,0,1,0,1,1,1,0,0,1],
     [1,1,0,1,1,0,1,1,0,1,1,1,0,0,0,1,0,1,1,0,1],
     [1,1,0,0,1,0,1,0,1,0,1,0,0,1,1,1,0,0,1,1,1],
     [1,0,1,0,0,1,1,1,1,1,0,1,0,1,0,1,0,0,1,0,1],
     [1,0,1,1,0,1,1,1,1,0,1,0,1,0,0,0,0,1,1,0,1],
     [1,0,1,1,1,1,1,0,0,0,0,1,0,0,1,1,1,1,1,1,1],
     [1,1,1,0,0,1,0,1,0,1,1,0,0,0,1,1,1,0,0,1,1],
     [1,1,0,1,1,1,1,0,0,1,0,0,1,1,0,1,1,0,0,1,0],
     [0,0,1,0,1,1,1,1,1,1,1,0,1,1,1,0,1,0,1,1,0],
     [0,1,1,1,1,1,0,1,0,0,0,0,1,1,1,1,0,1,0,1,0],
     [0,1,1,1,1,0,0,1,1,0,0,1,1,0,0,0,1,0,1,1,1],
     [0,0,1,0,1,0,1,0,1,1,1,1,1,0,0,1,0,1,0,1,1],
     [0,1,1,0,1,0,0,0,0,1,0,0,1,1,1,1,1,1,1,0,1],
     [1,1,1,0,0,0,0,1,1,1,1,1,1,1,0,0,0,1,1,1,0]]
@@
    Adjacency matrix of a 23-vertex graph with no $B_{2}$ in color 0 and no $B_{10}$ in color 1:
    [[0,0,1,1,1,1,1,1,0,1,1,1,1,1,1,0,0,1,0,0,1,1,0],
     [0,0,0,0,0,1,0,1,1,1,0,0,1,0,1,1,1,1,1,1,1,1,0],
     [1,0,0,1,1,0,1,0,0,0,1,1,0,1,1,1,1,1,0,0,0,1,1],
     [1,0,1,0,0,1,1,0,0,0,1,1,1,1,0,1,0,0,0,1,1,1,1],
     [1,0,1,0,0,0,1,1,1,1,1,1,0,1,1,0,1,1,1,0,0,0,1],
     [1,1,0,1,0,0,1,1,0,1,0,0,1,1,1,0,0,0,1,1,1,1,0],
     [1,0,1,1,1,1,0,1,0,0,0,1,0,1,1,0,0,0,1,1,0,1,1],
     [1,1,0,0,1,1,1,0,1,1,0,1,0,0,1,0,0,1,1,1,1,1,0],
     [0,1,0,0,1,0,0,1,0,1,1,1,1,0,0,1,1,1,1,1,1,0,1],
     [1,1,0,0,1,1,0,1,1,0,1,0,1,1,1,0,1,1,1,0,1,0,0],
     [1,0,1,1,1,0,0,0,1,1,0,1,1,1,0,1,1,1,0,0,1,0,1],
     [1,0,1,1,1,0,1,1,1,0,1,0,0,0,0,1,0,1,1,1,1,1,1],
     [1,1,0,1,0,1,0,0,1,1,1,0,0,1,0,1,1,0,1,1,1,0,1],
     [1,0,1,1,1,1,1,0,0,1,1,0,1,0,1,0,1,0,1,0,0,0,1],
     [1,1,1,0,1,1,1,1,0,1,0,0,0,1,0,0,1,1,1,0,0,1,0],
     [0,1,1,1,0,0,0,0,1,0,1,1,1,0,0,0,1,1,0,1,1,1,1],
     [0,1,1,0,1,0,0,0,1,1,1,0,1,1,1,1,0,1,1,0,0,1,1],
     [1,1,1,0,1,0,0,1,1,1,1,1,0,0,1,1,1,0,0,0,1,1,0],
     [0,1,0,0,1,1,1,1,1,1,0,1,1,1,1,0,1,0,0,1,0,0,1],
     [0,1,0,1,0,1,1,1,1,0,0,1,1,0,0,1,0,0,1,0,1,1,1],
     [1,1,0,1,0,1,0,1,1,1,1,1,1,0,0,1,0,1,0,1,0,1,0],
     [1,1,1,1,0,1,1,1,0,0,0,1,0,0,1,1,1,1,0,1,1,0,0],
     [0,0,1,1,1,0,1,0,1,0,1,1,1,1,0,1,1,0,1,1,0,0,0]]
@@
    Adjacency matrix of an 18-vertex graph with no $B_{3}$ in color 0 and no $B_{6}$ in color 1:
    [[0,1,1,1,1,0,0,1,1,0,1,1,0,0,0,0,1,1],
     [1,0,0,1,0,1,1,0,0,1,1,1,0,0,1,0,1,1],
     [1,0,0,1,0,0,1,1,1,1,1,0,0,1,0,1,0,1],
     [1,1,1,0,0,1,0,0,1,1,1,0,1,1,0,0,1,0],
     [1,0,0,0,0,1,0,1,1,1,0,1,1,0,0,1,1,1],
     [0,1,0,1,1,0,1,0,1,1,1,1,1,0,0,1,0,0],
     [0,1,1,0,0,1,0,0,1,1,0,1,0,1,1,1,0,1],
     [1,0,1,0,1,0,0,0,0,1,1,1,1,1,1,0,0,1],
     [1,0,1,1,1,1,1,0,0,0,0,1,0,1,0,1,1,0],
     [0,1,1,1,1,1,1,1,0,0,0,0,1,1,0,0,0,1],
     [1,1,1,1,0,1,0,1,0,0,0,1,1,0,1,1,0,0],
     [1,1,0,0,1,1,1,1,1,0,1,0,0,1,1,0,0,0],
     [0,0,0,1,1,1,0,1,0,1,1,0,0,1,1,1,1,0],
     [0,0,1,1,0,0,1,1,1,1,0,1,1,0,1,0,1,0],
     [0,1,0,0,0,0,1,1,0,0,1,1,1,1,0,1,1,1],
     [0,0,1,0,1,1,1,0,1,0,1,0,1,0,1,0,1,1],
     [1,1,0,1,1,0,0,0,1,0,0,0,1,1,1,1,0,1],
     [1,1,1,0,1,0,1,1,0,1,0,0,0,0,1,1,1,0]]
@@
    Adjacency matrix of a 19-vertex graph with no $B_{3}$ in color 0 and no $B_{7}$ in color 1:
    [[0,1,1,0,0,0,1,1,1,1,1,0,1,0,1,1,1,0,0],
     [1,0,1,0,1,0,1,1,1,1,0,1,0,0,0,0,1,1,1],
     [1,1,0,1,1,0,0,1,1,0,1,1,0,0,0,1,1,0,0],
     [0,0,1,0,0,1,1,0,1,0,1,1,1,1,0,0,1,1,0],
     [0,1,1,0,0,1,0,1,1,0,1,1,0,0,1,1,0,1,1],
     [0,0,0,1,1,0,1,1,1,1,1,0,0,1,1,0,1,1,0],
     [1,1,0,1,0,1,0,0,0,1,1,0,1,0,1,0,1,1,1],
     [1,1,1,0,1,1,0,0,0,1,1,0,0,1,0,1,1,1,0],
     [1,1,1,1,1,1,0,0,0,0,0,1,1,1,1,0,1,0,0],
     [1,1,0,0,0,1,1,1,0,0,0,0,0,1,1,1,1,0,1],
     [1,0,1,1,1,1,1,1,0,0,0,0,1,0,1,1,0,1,0],
     [0,1,1,1,1,0,0,0,1,0,0,0,1,1,0,1,0,1,1],
     [1,0,0,1,0,0,1,0,1,0,1,1,0,1,1,1,0,1,1],
     [0,0,0,1,0,1,0,1,1,1,0,1,1,0,0,1,1,1,1],
     [1,0,0,0,1,1,1,0,1,1,1,0,1,0,0,1,0,0,1],
     [1,0,1,0,1,0,0,1,0,1,1,1,1,1,1,0,0,0,1],
     [1,1,1,1,0,1,1,1,1,1,0,0,0,1,0,0,0,0,0],
     [0,1,0,1,1,1,1,1,0,0,1,1,1,1,0,0,0,0,1],
     [0,1,0,0,1,0,1,0,0,1,0,1,1,1,1,1,0,1,0]]
@@
    Adjacency matrix of an 18-vertex graph with no $B_{4}$ in color 0 and no $B_{5}$ in color 1:
    [[0,1,0,1,1,0,1,1,1,1,1,0,1,0,0,0,1,0],
     [1,0,0,0,1,1,1,0,0,1,0,0,0,1,1,0,1,1],
     [0,0,0,1,0,1,1,1,0,0,1,0,1,1,0,1,1,0],
     [1,0,1,0,0,0,1,1,1,0,0,1,0,1,1,0,1,0],
     [1,1,0,0,0,1,0,1,1,0,0,0,1,1,0,1,0,1],
     [0,1,1,0,1,0,0,0,1,1,1,1,0,1,0,1,1,0],
     [1,1,1,1,0,0,0,0,0,1,0,1,1,1,0,1,0,0],
     [1,0,1,1,1,0,0,0,0,1,1,0,0,1,1,1,0,1],
     [1,0,0,1,1,1,0,0,0,0,0,1,1,0,1,1,1,0],
     [1,1,0,0,0,1,1,1,0,0,1,1,0,0,1,1,0,0],
     [1,0,1,0,0,1,0,1,0,1,0,1,1,0,0,0,1,1],
     [0,0,0,1,0,1,1,0,1,1,1,0,1,1,1,0,0,1],
     [1,0,1,0,1,0,1,0,1,0,1,1,0,0,0,1,0,1],
     [0,1,1,1,1,1,1,1,0,0,0,1,0,0,0,0,0,1],
     [0,1,0,1,0,0,0,1,1,1,0,1,0,0,0,1,1,1],
     [0,0,1,0,1,1,1,1,1,1,0,0,1,0,1,0,0,0],
     [1,1,1,1,0,1,0,0,1,0,1,0,0,0,1,0,0,1],
     [0,1,0,0,1,0,0,1,0,0,1,1,1,1,1,0,1,0]]
@@
    Adjacency matrix of a 21-vertex graph with no $B_{4}$ in color 0 and no $B_{7}$ in color 1:
    [[0,0,0,1,0,0,1,0,0,0,1,1,1,1,1,1,1,0,1,0,0],
     [0,0,1,0,1,1,1,1,1,0,0,1,1,0,0,1,0,1,1,0,1],
     [0,1,0,1,0,0,1,1,0,1,1,1,0,0,0,1,0,1,0,1,1],
     [1,0,1,0,0,0,1,1,1,1,1,0,0,0,0,1,1,1,1,0,0],
     [0,1,0,0,0,1,1,0,1,1,0,0,1,0,1,0,1,0,1,1,1],
     [0,1,0,0,1,0,0,1,1,1,0,1,1,1,0,1,1,0,0,1,0],
     [1,1,1,1,1,0,0,0,0,0,0,1,0,0,1,1,1,0,1,1,1],
     [0,1,1,1,0,1,0,0,1,0,1,1,0,0,0,0,1,1,1,1,0],
     [0,1,0,1,1,1,0,1,0,1,0,1,0,1,1,0,0,1,1,0,0],
     [0,0,1,1,1,1,0,0,1,0,0,0,0,1,1,1,1,1,0,1,1],
     [1,0,1,1,0,0,0,1,0,0,0,1,1,1,1,0,0,1,0,1,1],
     [1,1,1,0,0,1,1,1,1,0,1,0,0,1,1,1,0,0,0,1,0],
     [1,1,0,0,1,1,0,0,0,0,1,0,0,1,0,1,1,1,1,0,1],
     [1,0,0,0,0,1,0,0,1,1,1,1,1,0,1,1,0,1,0,0,1],
     [1,0,0,0,1,0,1,0,1,1,1,1,0,1,0,0,0,0,1,1,1],
     [1,1,1,1,0,1,1,0,0,1,0,1,1,1,0,0,1,0,0,0,0],
     [1,0,0,1,1,1,1,1,0,1,0,0,1,0,0,1,0,0,1,1,0],
     [0,1,1,1,0,0,0,1,1,1,1,0,1,1,0,0,0,0,1,0,1],
     [1,1,0,1,1,0,1,1,1,0,0,0,1,0,1,0,1,1,0,0,0],
     [0,0,1,0,1,1,1,1,0,1,1,1,0,0,1,0,1,0,0,0,1],
     [0,1,1,0,1,0,1,0,0,1,1,0,1,1,1,0,0,1,0,1,0]]

### v2

@@ Abstract
In this work, we give several new upper and lower bounds on Ramsey numbers for books and wheels, including a tight upper bound establishing $R(W_{5},W_{7})=15$, matching upper and lower bounds giving $R(W_{5},W_{9})=18$, $R(B_{2},B_{8})=21$, and $R(B_{3},B_{7})=20$, and a number of additional tight lower bounds for books. We use a range of different methods: flag algebras, local search, bottom-up generation, and enumeration of polycirculant graphs.
@@ 1 Introduction
There are a number of results on Ramsey numbers for books and wheels, summarized in [rad_survey]. In particular, the first and third authors proved upper bounds for a variety of small Ramsey numbers, including for books and wheels, using the flag algebra method [LidP]; we establish most of the upper bounds here using the same method. The fourth author established a variety of lower bounds for graphs including wheels by explicit enumeration of circulant and polycirculant (or “block-circulant”) graphs [GOEDGEBEUR2022212]; here, we use the same technique to establish a number of new lower bounds for books. Other recent works have used reinforcement learning [RL_bounds] and SAT solvers [wesley] to give Ramsey lower bound constructions for books and wheels.
@@ 2 Results
Note: we have recently become aware of independent work establishing the same (tight) lower bound on $R(W_{5},W_{9})$ using an explicit construction [yanbo]. In [wesley2024], some of the results below on Ramsey numbers for books are also proven, and there is also a generalization of the second case of Lemma 1 to an infinite family.
@@
$R(W_{5},W_{7})$ 15* 15 [wesley] 16 [LidP]
$R(W_{5},W_{9})$ 18* 18* 17 [rad_survey]
$R(B_{2},B_{8})$ 21* 21* 19 [RS78] 22 [b2_uppers]
@@
$R(B_{2},B_{10})$ 25 23 [RS78] 26 [b2_uppers]
$R(B_{3},B_{7})$ $\textit{20}^{\star}$ $\textit{20}^{\star}$
$R(B_{4},B_{7})$ 22 23
@@
We also have the following lemma covering the diagonal and almost-diagonal cases for some additional sizes of books. Note that some of these bounds are also included in the table above for readers’ convenience.
Lemma 1.
For all $4\leq n\leq 21$ we have
  • $4n-3\leq R(B_{n-2},B_{n})$
  • $R(B_{n-1},B_{n})=4n-1$
  • $24n+1\leq R(B_{n},B_{n})\leq 4n+$ 2
Note that it was also already known that $R(B_{n-2},B_{n})\leq 4n-3$ if $n\equiv 2\pmod{3}$, and that $R(B_{n},B_{n})\leq 4n-1$ if $14n+$ 1 is not the sum of two squares.
@@ 3 Method
Each of the lower bounds in the table above is established by a construction provided in Appendix A. The constructions were found using tabu search or as polycirculant graphs, and verified independently in SageMath [sagemath]; the code for books and wheels is available at https://github.com/gwen-mckinley/ramsey-books-wheels. The additional lower bounds given in Lemma 1 were also established by polycirculant graph constructions, which are available in the same repository. The majority of the upper bounds were found by semidefinite programming using the flag algebra method, as described in [LidP]. The certificates for the upper bounds are available at http://lidicky.name/pub/gr. And several upper bounds are from “bottom-up generation,” as described below.
Here is more detail on the structure of some lower bound constructions for generalized Ramsey numbers: for $GR(3,K_{4},2)$, one extremal coloring is composed of the Paley graph on 9 vertices in the first color, and the other two colors are 3 triangles each. This is the only coloring with more than 6 symmetries. For $GR(4,K_{4},3)$, the only maximal graph is isomorphic to $3K_{3}$ in every color. Joining two arbitrary colors, we find the graph described for $GR(3,K_{4},2)$. One example for $GR(3,K_{5},2)$ is the coloring composed of the (isomorphic) circulant graphs $C(19,[1,7,8])$, $C(19,[2,3,5])$ and $C(19,[4,6,9])$.
3.1 Bottom-up generation
For small Ramsey numbers, it is possible to generate all Ramsey graphs by starting from a trivial case (for example the graph on one vertex) and repeatedly adding an extra vertex in every possible way without creating a subgraph isomorphic to one of the avoided graphs. For two-color Ramsey numbers this can be achieved efficiently with the program geng from the nauty package written by Brendan McKay [MCKAY201494]. We wrote a custom plugin for generating Ramsey graphs avoiding wheels and books. This allowed us to find the values of $R(W_{5},W_{7})$ and $R(W_{5},W_{9})$ in 90s and 36 CPU-hours respectively. The computation of $R(B_{2},B_{8})$ was significantly heavier and required approximately 2 years of CPU time. The counts are presented in Table 2. To test the latter program, we also generated all $R(B_{2},B_{6})$-graphs (which requires only a trivial modification to the program) and verified that the results were identical to those in [BlackLevenRadz].
For the generalized Ramsey numbers, multiple colors are needed, which is not directly supported by nauty. We therefore used a simple implementation in SageMath [sagemath] to do the generation. Both $GR(3,K_{4},2)$ and $GR(4,K_{4},3)$ could be solved in less than one minute in this way.
3.2 Flag Algebras
The new upper bounds (except for $R(B_{2},B_{8})$) were obtained (independently) using the flag algebra method. This method was developed by Razborov [MR2371204], and the main idea is to find a certificate for the bound as a sum of squares, which can be done efficiently using semidefinite programming. We used CSDP [csdp] and MOSEK [mosek] solvers in this work. While the solvers produce numerical results, we created tools to change the solution to an exact arithmetic one in SageMath [sagemath]. The flag algebra method is quite general, and it is naturally applicable to problems where one is interested in asymptotic results. For example, the Turán density of small hypergraphs and its variants [l2norm] is a group of problems where flag algebras led to many new improvements.
3.3 Polycirculant graphs
$k$-polycirculant graphs are graphs of order $n$ that have a non-trivial automorphism $\alpha$ such that all vertex orbits under $\alpha$ have the same size $n/k$. For a large amount of Ramsey numbers, there exist extremal graphs that are polycirculant. One might also believe that if an extremal graph is unique, it is more likely to have some symmetries. For example, the unique Ramsey-$(B_{2},B_{8})$-graph on 20 vertices has 240 automorphisms and is 4-polycirculant.
We used the exhaustive algorithm outlined in [GOEDGEBEUR2022212] and adapted it slightly to generate all polycirculant Ramsey graphs avoiding books. With this generator, it is for example easy to show that there are exactly seven 2-polycirculant Ramsey-$(B_{2},B_{9})$-graphs on 20 vertices and exactly one 3-polycirculant Ramsey-$(B_{2},B_{10})$-graph on 24 vertices (up to isomorphism).
Proof of Lemma 1.
The upper bounds were proven in [RS78]. The lower bounds are witnessed by 2-polycirculant graphs (as was in fact one of the examples in [RS78]), which are available here: https://github.com/gwen-mckinley/ramsey-books-wheels
Note: the polycirculant graph constructions we found all have the extra property that the first circulant block (the subgraph induced by the first vertex-orbit of the 2-polycirculant automorphism) is isomorphic to the complement of the second circulant block. For most cases there were many 2-polycirculant Ramsey graphs with this property. For example: there are 581 non-isomorphic 2-polycirculant $(B_{12},B_{13})$-graphs on 50 vertices. Most of these graphs had no extra automorphisms besides the ones implied by being 2-polycirculant. This makes it hard to deduce a pattern in these graphs that could generalize to an infinite family.
3.4 Tabu search
3.4.1 Tabu search details
@@ 4.1 Upper bounds
4.2 Lower bounds
4.2.1 Poycirculant graphs
It is interesting to ask whether Lemma 1 can be extended to larger values of $n$. We proved this lemma by starting a new search for polycirculant graphs for each $n$. There was no indication that the pattern would stop at $n\geq 22$; the generation simply started to take a lot of time. It seems plausible that there is some unified family of graphs giving this bound for all $n$. But again, as noted in Section 3.3, most of the graphs have no additional symmetry beyond being polycirculant, making it difficult to discern any larger pattern. Note that a partial answer to this question was given in [wesley2024] where the lemma is generalized to an infinite family consisting of orders that are twice a prime power (that is 1 modulo 4).
4.2.2 Tabu search
To extend the tabu search approach used in this work for finding lower bounds, there are a number of natural directions to explore. Perhaps most obvious would be to implement everything in a lower-level language – all the local search code for wheels and books is currently written in Python. In particular, the score-change functions dominate the total cost, and speeding them up would almost certainly improve performance; relevant here is an extensive literature on subgraph counting (see [subgraph_survey]), and it may be possible to simply incorporate existing open-source solutions.
Last but not least, one could try to be more strategic in choosing where to start a local search. One common approach is to build up from smaller graphs: we first find a construction on a smaller number of vertices avoiding “bad” subgraphs, then add a vertex, and run our local search starting from this new graph (as opposed to starting from a random graph) [whatsR55, exoo4_6]. The hope is that by doing this, we direct the search to a promising area of the (very large) search space. We also tried this method, with limited success, but it is likely worth further exploration.
@@ Acknowledgments
The second author would like to thank David Hemminger and Larry Wilson for many helpful and stimulating discussions, and for their consistent encouragement. This work used the computing resources at the Center for Computational Mathematics, University of Colorado Denver, including the Alderaan cluster, supported by the National Science Foundation award OAC-2019089. We also used the Stevin computing infrastructure provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation Flanders (FWO) and the Flemish Government.
@@
Appendix A Counts and constructions
$n$ $R(B_{2},B_{8})$ $R(W_{5},W_{7})$ $R(W_{5},W_{9})$ $GR(3,K_{4},2)$ $GR(4,K_{4},3)$
1 1 1 1 1 1
2 2 2 2 1 1
3 4 4 4 3 3
4 9 11 11 9 7
5 22 31 31 34 11
6 69 130 130 154 12
7 255 675 723 428 1
8 1301 4868 6456 556 1
9 9297 38059 87977 263 1
10 96618 251377 1627532 0 0
11 1405777 878658 27891376
12 25330324 932411 250459368
13 443322144 141871 509767930
14 5130980404 1161 139131233
15 ? 0 46736023
16 ? 18956337
17 ? 272891
18 ? 0
19 41
20 1
21 0
Table 2: Counts for Ramsey graph. The multicolor numbers are given up to color-swapping isomorphism. Because the calculations for $R(B_{2},B_{8})$ had to be split up into multiple processes, we have no accurate counts for some orders.
@@ Appendix A Counts and constructions
    Graph6-string of a 14-vertex graph with no $W_{5}$ as a subgraph and no $W_{7}$ in the complement:
    Mav?Hwu]`ySZpyyg?
@@
    Graph6-string of a 17-vertex graph with no $W_{5}$ as a subgraph and no $W_{9}$ in the complement:
    PIL{eMI^Jqp[gXkp_|zxOaww
@@
    Graph6-string of a 20-vertex graph with no $B_{2}$ as a subgraph and no $B_{8}$ in the complement:
    SXgcISrdSaJQBJs_jp@CWFOV?q}HWOPbc
@@
    Graph6-string of a 21-vertex graph with no $B_{2}$ as a subgraph and no $B_{9}$ in the complement:
    TXhJ?ScLQoHAO]EhcLe_G_nAEXuiBSnW?]?w
@@
    Graph6-string of a 24-vertex graph with no $B_{2}$ as a subgraph and no $B_{10}$ in the complement:
    W?bFFbw^@{BwgDsAl?lg@U_cl@GlDGUacDih?lTKApSgDqh
@@
    Graph6-string of a 18-vertex graph with no $B_{3}$ as a subgraph and no $B_{6}$ in the complement:
    QG]cql@_GTeAwhrAVeEaiHv?Sn?
@@
    Graph6-string of a 19-vertex graph with no $B_{3}$ as a subgraph and no $B_{7}$ in the complement:
    REf`OcBMI@ozZSyaMGil?ABm_o|jOO
@@
    Graph6-string of a 18-vertex graph with no $B_{4}$ as a subgraph and no $B_{5}$ in the complement:
[…]
@@
    Graph6-string of a 21-vertex graph with no $B_{4}$ as a subgraph and no $B_{7}$ in the complement:
[…]